OpenStudy (anonymous):

For the school concert band performance, student tickets cost $8 each and adult tickets cost $12 each. The sellers collected $3,200 from 300 tickets sold. How many adult tickets and student tickets were sold?

OpenStudy (kirbykirby):

Let \(a: \text{number of adult tickets}\) \(s: \text{number of student tickets}\) Since there are 300 tickets sold, you can say that you have: \(a + s = 300\) And since the student ticket costs $8, you can say that the total amount of student tickets to be \(8s\) And the total amount of adult tickets to be \(12a\) So, in total, they sold $3200 in tickets, so: \(8s+12a=3200\) So you have 2 equations and 2 unknowns. You can now solve the system of equations by whichever method you prefer. \(a + s = 300\) \(8s+12a=3200\)